OpenStudy (anonymous):

If the sum of 2 nos. is s find the minimum value of the sum of their squares..

OpenStudy (anonymous):

nos?

OpenStudy (anonymous):

numbers

OpenStudy (anonymous):

x+y=s minimize x^2+y^2? graph

OpenStudy (anonymous):

x^2+y^2 = (x+y)^2 -2xy .... maximum value of this occurs when one of the x or y is zero and other is six.... In this case, x^2+y^2 = (x+y)^2 = s^2

OpenStudy (anonymous):

i dont u are allowed to use calculus or not...so i use another method we have\[x+y=s\]so\[x^2+y^2=x^2+(s-x)^2=2x^2-2sx+s^2\]this is an upward quadratic since (leading coefficient is positive) so the maximim value is \(\infty\) but it would be a nice try if u find minimum value of \(x^2+y^2\) if \(x+y=s\)

OpenStudy (anonymous):

*i dont know

OpenStudy (anonymous):

oh the problem is finding minimum...sorry...so u just need to minimize\[y=2x^2-2sx+s^2\]

OpenStudy (anonymous):

Just use that in axx+bx+c=y, the minimum point on the graph is when x=-b/2a